# Recovery of Quantum Correlations using Machine Learning

### v2

@@ I Introduction
Recent advances have demonstrated the utility of ML techniques in quantum information science, such as continuous-variable quantum key distribution (CV-QKD) [15], quantum error correction [16], quantum control [17], and particularly in quantum state tomography (QST) [18], which aims to reconstruct the complete quantum state of a system, represented by its density matrix, based on experimental measurements. This process is crucial for verifying quantum system behavior, characterizing quantum devices, analyzing quantum phenomena such as entanglement and coherence, and benchmarking experimental results against theoretical predictions [19, 20]. Machine-learning-based QST (ML-QST) has emerged as a transformative approach to quantum state reconstruction, addressing key challenges of scalability and noise in traditional QST [21]. Leveraging advanced ML techniques such as convolutional neural networks [22, 23], adaptive neural networks [24], deep learning [25], and most recently quantum machine learning [18], ML-QST enables efficient processing of large datasets, reduces measurement requirements, and enhances noise resilience [26]. It is versatile, applicable across diverse quantum platforms, and effective at handling high-dimensional systems [24]. Unlike ML-QST, which reconstructs a complete steady state of a quantum system, our approach focuses on recovering dynamic quantum correlations within a quantum system, rather than reconstructing the quantum state itself, based on partial or indirect information. Specifically tailored for dynamic scattering scenarios, our method employs time-series predictions to restore quantum correlations in a quantum system, offering a distinct and complementary perspective.
@@ V Conclusion and Discussion
Distinction from the hardware method. The hardware method [13] physically recovers quantum correlations by recollecting scattered photons, directly mitigating photon loss. In contrast, our ML-based approach reconstructs disrupted quantum correlations in a quantum system (rather than the intact state of the quantum system in ML-aided QST) based on prior correlations and system characteristics. This distinction indeed parallels the difference between quantum error correction (which protects real quantum states) and quantum error mitigation (which focuses on correcting measurement outcomes). While it is true that our software-based approach may not directly address quantum communication tasks, which require transmission of physical quantum states, its utility should not be underestimated for other applications. Specifically, the software-based method offers potential for certain pre-transmission tasks, such as noise mitigation and preparation of quantum resources, even if not suitable for direct quantum state transmission. Also, the software-based method eliminates the need to localize the scatterer, simplifying the experimental setup. This could be particularly advantageous in scenarios where environmental disturbances are unpredictable or cannot be directly addressed through hardware. Rather than being viewed as mutually exclusive, the hardware and software methods can complement each other. For example, hardware approaches may handle primary recovery of lost photons, while software methods further enhance system resilience by mitigating residual decoherence effects.
Potential applications. While our software approach does not physically recover lost photons, it serves as a powerful tool to infer and reconstruct quantum correlations in a quantum system in scenarios where physical recovery is either impractical or impossible. For example, in situations with severe scattering or inaccessible environments, our ML-based method provides a means to restore useful quantum correlations without hardware modifications. Specifically, the ML-based method can be employed in: 1) Scenarios with high loss: where physical recovery of photons is infeasible, such as in distributed quantum networks or highly scattering environments; 2) Measurement error mitigation: by extrapolating missing information based on known system dynamics, the method can enhance the robustness of quantum systems for tasks like sensing, metrology, or pre-transmission signal preparation; and 3) Cost-effective solutions: the software-based approach reduces the need for complex hardware modifications, making it more accessible for certain applications.

### v3

@@ I Introduction
Recent advances have demonstrated the utility of ML techniques in quantum information science, such as continuous-variable quantum key distribution (CV-QKD) [15], quantum error correction [16], quantum control [17], and particularly in quantum state tomography (QST) [18], which aims to reconstruct the complete quantum state of a system, represented by its density matrix, based on experimental measurements. This process is crucial for verifying quantum system behavior, characterizing quantum devices, analyzing quantum phenomena such as entanglement and coherence, and benchmarking experimental results against theoretical predictions [19, 20]. Machine-learning-based QST (ML-QST) has emerged as a transformative approach to quantum state reconstruction, addressing key challenges of scalability and noise in traditional QST [21]. Leveraging advanced ML techniques such as convolutional neural networks [22, 23], adaptive neural networks [24], deep learning [25], and most recently quantum machine learning [18], ML-QST enables efficient processing of large datasets, reduces measurement requirements, and enhances noise resilience [26]. It is versatile, applicable across diverse quantum platforms, and effective at handling high-dimensional systems [24]. Unlike ML-QST, which reconstructs a complete static state of a quantum system, our approach focuses on recovering dynamic quantum correlations within a quantum system, rather than reconstructing the quantum state itself, based on partial or indirect information. Specifically tailored for evolving scattering scenarios, our method employs time-series predictions to restore quantum correlations in a quantum system, offering a distinct and complementary perspective, which also differentiates it from static beamsplitter-modeled analytical loss formulas.
@@ V Conclusion and Discussion
Distinction from the hardware method. The hardware method [13] physically recovers quantum correlations by recollecting scattered photons, directly mitigating photon loss. In contrast, our ML-based approach reconstructs disrupted evolving quantum correlations in a quantum system (rather than the full static state reconstruction in ML-aided QST) based on prior correlations and system characteristics. The distinction between the hardware and software methods indeed resemblances the difference between quantum error correction (which protects real quantum states) and quantum error mitigation (which focuses on correcting measurement outcomes). While it is true that our software-based approach may not directly address quantum communication tasks, which require transmission of physical quantum states with high fidelity, its utility should not be underestimated for other applications. Specifically, the software-based method eliminates the need to localize the scatterer, which is particularly advantageous in scenarios where environmental disturbances are unpredictable or cannot be directly addressed through hardware approach, such as in atmospheric quantum communication. Rather than being viewed as mutually exclusive, the hardware and software methods can complement each other. For example, hardware approaches may handle primary recovery of lost photons, while software methods further enhance system resilience by mitigating residual decoherence effects.
Potential applications. While our software approach does not physically recover lost photons, it serves as a powerful tool to infer and reconstruct quantum correlations in a quantum system in scenarios where physical recovery is either impractical or impossible. For example, in situations with severe scattering or inaccessible environments, our ML-based method provides a means to restore useful quantum correlations without hardware modifications. Specifically, the ML-based method can be employed in: (1) Scenarios with high loss: where physical recovery of photons is infeasible, such as in distributed quantum networks or highly scattering environments (e.g., atmospheric quantum communication); (2) Measurement error mitigation: by extrapolating missing information based on known system dynamics, the method can enhance the robustness of quantum systems for tasks like sensing or pre-transmission signal preparation; and (3) Cost-effective solutions: the software-based approach reduces the need for complex hardware modifications, making it an accessible alternative for resource-limited settings.
It is worth noting that our LSTM-based ML approach can extend dynamical decoupling (DD), originally designed to mitigate decoherence and dissipation in open quantum systems [47], to quantum communication by predicting optimal pulse sequences. By analyzing real-time environmental noise and quantum correlation evolution, the LSTM model can dynamically adjust DD pulse timing and strength, potentially enhancing the performance of standard sequences like Carr-Purcell-Meiboom-Gill and Uhrig [48]. This adaptive control can mitigate qubit dephasing [49] and entanglement decay [50] in fiber-based quantum communication, providing software-driven noise suppression that improves reliability and optimizes resource efficiency without additional hardware modifications.